[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue May 31, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue May 31 14:12:15 UTC 2016


HSU Shift Log for Tue May 31 2016 10:12:10 GMT-0400 (EDT)

Shift: H

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: There was an issue with sending and retrieving the 12Z  
guidance for AL922016, possibly due to the re-numbering from AL022016.  
However, the guidance was successfully retrieved.

N-AWIPS/AWIPS2 issues: None.

Other Techincal issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160531/0649718b/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list